OpenStudy (anonymous):

help For a regular n-gon: a. What is the sum of the measures of its angles? b. What is the measure of each angle? c. What is the sum of the measures of its exterior angles, one at each vertex? d. What is the measure of each exterior angle? e. Find the sum of your answers to parts b and d. Explain why this sum makes sense.

OpenStudy (anonymous):

I tried answering this but I was lost

OpenStudy (anonymous):

sum of exterior angles of a polygon=360 if it is regular each exterior angle=360/n each interior angle=180-360/n \[=180\left( \frac{ n-2 }{n } \right)\] sum of interior angles=multiply by n